Julie performs an experiment in which she rolls 5 fair 6-sided dice and records the number XX of dice that have a value of at least 3. For example, if the rolls are {2, 4, 1, 6, 4}, then X=3
a) The random variable X follows a binomial distribution. What are n and p for this distribution?
n=Answer
(Enter a decimal only, not a fraction, rounded to the nearest hundredth, in the box below.)
p=Answer

b) What is the probability that exactly 1 of the 5 dice has a value of at least 3?
(Enter a decimal only, not a fraction, rounded to the nearest hundredth, in the box below.)

Answer by Boreal(15235)

n=5
p=2/3 or 0.67, since 3,4,5,6 work and 1,2 do not.
================
exactly 1 has a value of at least 3.
5 ways it can happen * (2/3)^1(1/3)^4=5*2/243=10/243 or 0.0412. Note, if fractions aren't used, the number is not as accurate and would be 0.0397